4)If 10 persons are to sit in a round table then the odds against two specific persons sitting together is
a)1/9 b) 2/9 c)7 to 2 d)2 to 7

Grade:12

1 Answers

Nishant Vora IIT Patna
7 years ago
Total number of ways = 9!

favorable ways = 8! 2!

Hence P =8! 2! /9! = 2/9 = 2/ 2+7

hence odd agains are 7 : 2